OpenStudy (anonymous):

Find the derivative f(x)=3x^4-9x^3+5

jimthompson5910 (jim_thompson5910):

hint: use the power rule if y = x^n, then y ' = n*x^(n-1)

OpenStudy (anonymous):

yeah thats what i was thinking which would make it =12x^3-27x^2 but do I also use the product rule?

jimthompson5910 (jim_thompson5910):

no need for the product rule since you don't have a product of two variable terms

jimthompson5910 (jim_thompson5910):

what you have is the correct final answer

OpenStudy (anonymous):

thats it?? nothing else to do? and you drop the constant since its derivative is zero right?

jimthompson5910 (jim_thompson5910):

you are correct, the derivative of a constant function is 0

jimthompson5910 (jim_thompson5910):

and yes, nothing else can be done

OpenStudy (anonymous):

ok guess i am trying to make this harder than it is Thanks

jimthompson5910 (jim_thompson5910):

you're welcome